# The CLC-UKET Dataset: Benchmarking Case Outcome Prediction for the UK Employment Tribunal

### v1

@@ 7.2 Possible Reasons for Errors
Models and annotators, based on the extracted facts and claims, cannot always determine whether a tribunal’s decision will finally resolve the claim or only address a preliminary issue. For example, in a disability discrimination case, the tribunal might first issue a judgment confirming the claimant’s disability (preliminary issue), followed by a second judgment addressing the actual discrimination claim. The first judgment (which the claimant may win) is a necessary step but does not resolve the final claim, while the second judgment might conclude that there was no discrimination (such that the claimant ultimately loses). Preliminary issues are often contested, and some applications may solely seek a tribunal declaration on such issues (e.g., confirming the claimant is an employee or disabled). The possibility of such multi-step proceedings increases the complexity of predictions and has likely had a negative effect on the results for both the models and the human predictors.
Further difficulties arise in cases where UKET renders a procedural decision instead of deciding on the substance of the claim. Such cases are classified as “other”. However, both models and human annotators may predict a substantive instead of a procedural decision and, therefore, suggest “claimant wins” or “claimant loses”. According to our annotation guidelines, this affects, in particular, the categories of “claimant partly wins” and “other”. This complexity may have contributed to low evaluation scores for “claimant partly wins” and “other”.
@@ 8 Conclusion
This paper explores the prediction of dispute outcomes for the UK Employment Tribunal (UKET). The paper demonstrates efforts on employing LLMs for automatic annotation practice to alleviate the burden of massive manual annotation. With LLM-aided annotation, we curated the CLC-UKET dataset with comprehensive, high-quality legal annotations. We showcased how the CLC-UKET data can be used to construct a prediction task to categorise case outcomes based on sequences of facts and claims. We fine-tuned and evaluated two widely used Transformer-based models on this prediction task. We also evaluated LLMs on the prediction task with a range of settings, and reported human performance on the task to facilitate model calibration. These empirical efforts serve as a useful benchmark for the UKET prediction task. We will make the CLC-UKET dataset publicly available with the publication of this paper.
@@
The curated dataset is developed on the basis of the Cambridge Law Corpus (CLC), which aggregates publicly available UK legal judgments. Both the decisions in the CLC and the jurisdiction codes of UKET are licensed for use under the Open Government Licence. This licence grants a worldwide, royalty-free, perpetual and non-exclusive licence. Access to the CLC is restricted to researchers with confirmed ethical clearance and requires compliance with the DPA and UK GDPR. While UK legal judgments are not anonymised, Rule 50 of the Employment Tribunal Rules ensures that sensitive personal information is anonymised when necessary. Additionally, Schedule 2, Part 5 of the DPA provides derogations for academic research, alleviating the burden of notifying all individuals involved in judgments.
@@
While our study provides valuable insights into the prediction of dispute outcomes for the UK Employment Tribunal, it is important to acknowledge certain limitations of our findings.
@@
Automatic information extraction. Manual annotation of legal texts requires extensive expert knowledge and can be costly. To alleviate these challenges, this research utilised GPT-4 for automatic information extraction. While the use of GPT-4 offers notable advantages in terms of time and cost efficiency, and the extraction results are generally satisfactory according to the quality check conducted by legal experts in a related study (de Faria et al., 2024), this annotation practice is not without flaws. The quality of legal annotations could be further improved in future explorations. There is also room to explore the effect of extracting and providing more detailed facts compared to the relatively concise fact statements present in the current CLC-UKET dataset.
@@ A.1 Experiment Settings for Transformer Models
The implementation of the two Transformer-based models is based on the HuggingFace transformer library (Wolf et al., 2020). We used the base versions for both models, initialised from their pre-trained weights. The BERT-base checkpoint has 110 million parameters. The T5-base checkpoint has 220 million parameters. The maximum input sequence length was set to 512 tokens. We tried different settings for other hyperparameters such as weight decay and the number of warm-up steps, and found that the values of those hyperparameters have an impact on how fast the model is trained, especially at the beginning steps, but do not have a strong impact on the final learning performance. For this reason, we set both weight decay and warm-up steps to 0 for ease of model implementation and future replication. All training processes were performed on an Nvidia RTX 8000 GPU.
@@ A.2 Final Prompts Used in the GPT-based Experiments
You are a legal assistant. Your task is to read through the court decisions that I will send you, and extract the following information for each input: 1. facts of the case; 2. claims made; 3. any references to legal statutes, acts, regulations, provisions and rules, including the specific number(s), section(s) and article(s) of each of them, and including procedural tribunal rules; 4. references to precedents and other court decisions; 5. general case outcome; 6. general case outcome summarised using one of the four labels - ‘claimant wins’, ‘claimant loses’, ‘claimant partly wins’ and ‘other’; 7. detailed order and remedies; 8. essential reasons for the decision (procedural and substantial). If there are multiple claimants or respondents, extract the case outcome for each and all of the claimants or respondents separately. Please stick strictly to the text contents that I will send.
@@ Appendix C Aggregating “wins” and “partly wins”
Appendix D Annotation Guidelines for Human Prediction of UKET Case Outcomes
D.1 Introduction

### v2

@@ 7.2 Possible Reasons for Errors
Models and annotators, based on the extracted facts and claims, cannot always determine whether a tribunal’s decision will finally resolve the claim or only address a preliminary issue. For example, in a disability discrimination case, the tribunal might first issue a judgment confirming the claimant’s disability (preliminary issue), followed by a second judgment addressing the actual discrimination claim. The first judgment (which the claimant may win) is a necessary step but does not resolve the final claim, whilst the second judgment might conclude that there was no discrimination (such that the claimant ultimately loses). Preliminary issues are often contested, and some applications may solely seek a tribunal declaration on such issues (e.g., confirming the claimant is an employee or disabled). The possibility of such multi-step proceedings increases the complexity of predictions and has likely had a negative effect on the scores of both the models and the human predictors.
Further difficulties arise in cases where the UKET renders a procedural decision instead of deciding on the substance of the claim. Such cases are classified as “other”. However, both models and human annotators may predict a substantive instead of a procedural decision and, therefore, suggest “claimant wins” or “claimant loses”. According to our annotation guidelines, this affects, in particular, the categories of “claimant partly wins” and “other”. This complexity may have contributed to low evaluation scores for “claimant partly wins” and “other”.
@@ 8 Conclusion
This paper explores the prediction of dispute outcomes for the UK Employment Tribunal (UKET). It also illustrates the utility of LLMs for automatic annotation to reduce the burden of extensive manual annotation. With LLM-aided annotation, we curated the CLC-UKET dataset with comprehensive, high-quality legal annotations. We showcased how the CLC-UKET data can be used to construct a prediction task to categorise case outcomes based on sequences of facts and claims. We fine-tuned and evaluated two widely used Transformer-based models on this prediction task. We also evaluated LLMs on the prediction task with a range of settings, and reported human performance on the task to facilitate model calibration. These empirical efforts serve as a useful benchmark for the UKET prediction task. We will make the CLC-UKET dataset publicly available [^8] The CLC website: https://www.cst.cam.ac.uk/research/srg/
projects/law.
to facilitate future research in this field.
@@
The curated dataset is developed on the basis of the Cambridge Law Corpus (CLC), which aggregates publicly available UK legal judgments. Both the decisions in the CLC and the jurisdiction codes of UKET are licensed for use under the Open Government Licence. This licence grants a worldwide, royalty-free, perpetual and non-exclusive licence. Access to the CLC is restricted to researchers with confirmed ethical clearance and requires compliance with the DPA and UK GDPR. Whilst UK legal judgments are not anonymised, Rule 50 of the Employment Tribunal Rules ensures that sensitive personal information is anonymised when necessary. Additionally, Schedule 2, Part 5 of the DPA provides derogations for academic research, alleviating the burden of notifying all individuals involved in judgments.
@@
Whilst our study provides valuable insights into the prediction of dispute outcomes for the UK Employment Tribunal, it is important to acknowledge certain limitations of our findings.
@@
Automatic information extraction. Manual annotation of legal texts requires extensive expert knowledge and can be costly. To alleviate these challenges, this research utilised GPT-4 for automatic information extraction. Whilst the use of GPT-4 offers notable advantages in terms of time and cost efficiency, and the extraction results are generally satisfactory according to the quality check conducted by legal experts in a related study (de Faria et al., 2024), this annotation practice is not without flaws. The quality of legal annotations could be further improved in future explorations. There is also room to explore the effect of extracting and providing more detailed facts compared to the relatively concise fact statements present in the current CLC-UKET dataset.
@@ A.1 Experiment Settings for Transformer Models
The implementation of the two Transformer-based models is based on the HuggingFace transformer library (Wolf et al., 2020). We used the base versions for both models, initialised from their pre-trained weights. The BERT-base checkpoint has 110 million parameters. The T5-base checkpoint has 220 million parameters. The maximum input sequence length was set to 512 tokens [^9] All input texts to BERT and T5 are under this token length limit.. We tried different settings for other hyperparameters such as weight decay and the number of warm-up steps, and found that the values of those hyperparameters have an impact on how fast the model is trained, especially at the beginning steps, but do not have a strong impact on the final learning performance. For this reason, we set both weight decay and warm-up steps to 0 for ease of model implementation and future replication. All training processes were performed on an Nvidia RTX 8000 GPU.
@@ A.2 Final Prompts Used in the GPT-based Experiments
You are a legal assistant. Your task is to read through the court decisions that I will send you, and extract the following information for each input: 1. facts of the case; 2. claims made; 3. any references to legal statutes, acts, regulations, provisions and rules, including the specific number(s), section(s) and article(s) of each of them, and including procedural tribunal rules; 4. references to precedents and other court decisions; 5. general case outcome; 6. general case outcome summarised using one of the four labels - ‘claimant wins’, ‘claimant loses’, ‘claimant partly wins’ and ‘other’; 7. detailed order and remedies; 8. essential reasons for the decision (procedural and substantive). If there are multiple claimants or respondents, extract the case outcome for each and all of the claimants or respondents separately. Please stick strictly to the text contents that I will send.
@@ Appendix C Aggregating “wins” and “partly wins”
Appendix D Human Prediction for UKET Case Outcomes
D.1 Annotation guidelines
D.1.1 Introduction
@@ D.1.3 Annotation instructions
Label Kappa Score
claimant wins 0.322
claimant loses 0.191
claimant partly wins 0.284
other 0.470
Table 9: Annotators’ agreement across four label categories, measured by Cohen’s Kappa scores.
Appendix E Annotators’ agreement across label categories
In Table 9 we report Cohen’s Kappa scores for the predictions of two annotators under four label categories - claimant wins, claimant loses, claimant partly wins and other.